OpenStudy (anonymous):

What are the solutions of the quadratic equation below? -7x2 - 23x + 10 = 0

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

\[-7x2 - 23x + 10 = 0\] right?

OpenStudy (misty1212):

probably best to rewrite it as \[7x^2+ 23x-10 = 0\]

OpenStudy (anonymous):

they are not here @wolf1728 just wasting your time

OpenStudy (wolf1728):

x = [-b +-sq root(b^2 -4*a*c)] / 2*a x = [-23 +-sq root(529 -280)] / 14 x = [-23 +-(15.7797338381)] / 14

OpenStudy (wolf1728):

jasdeleon - it seems you are back